RUN 2B L1CAL MEETING MINUTES 16 October, 2003 Present o Fermilab: D.Calvet, G.Cancelo, P.Laurens, J.Mitrevski, V.Pavlicek, E.Perez, V.O'Dell, S.Rapisarda o MSU: M.Abolins o Nevis: H.Evans o Rice: P.Padley Integration Tests: ------------------ o Integration Test has been a huge success so far! - thanks to Denis, Emmanuelle, Dan, Philippe, Jovan and John o Goals Acccomplished 1) ADF gets timing from SCL via ADF Timing Test Card 2) TAB gets timing from SCL via VME/SCL card 3) Data transfered from ADF to TAB . works both with ADC's on and off . only ~32 events at a time read out and verified . no sync between start of ADF data and start of TAB > requires SCL Init to be seen at ADF . temporary connection of cable to TAB is tricky > will be improved with full TAB backplane 4) TAB to Cal-Track (actually L1Muon) data transfer . note: this was not in the original plan - but things were going so well that we tried it anyway . timing of sync-gap signals checked on scope > this has historically been the hard part . no mechanism to dump data at L1Muon end, but... . have sent known patterns (1 e per 32 events, 2 e per 32 events, etc.) and seen consistent changes in L1Muon trigger rates > also have checked that these triggers occur in correct timing bin 5) TWG data transfered to ADF . see below 6) Large volumes of data between ADF and TAB . check for errors using parity . goal is to produce estimate of bit error rate . firmware to calculate parity debugged . no parity errors observed on serializers 1 and 2 > constant data pattern > run for ~1 minute . some errors on serializer 3 > may be some noise on this o Still on the List 1) Continue large data volume tests 2) TAB to L2 data transfer . note: note in original plan . if this can be verified then TAB is ready for production > after TAB-to-GAB data transfer tests at Nevis using TAB-to-GAB test card . will talk with Reinhardt about this 3) Test that can receive L1Accept, SCL Init on ADF Timing card Test Waveform Generator: ------------------------ o October 15 Tests with ADF - TWG, computer, scope brought to Test Area - sent sequence of short pulses to ADF o ADF Results - input signal 2V P-to-P - use cable harness for input of signals . 1st test of this with "real" signals . note two of the cables were broken - Channels 0-15 . ~7 working well, rest have some noise - Channels 16-31 . more noise observed - consistent with measurements at Saclay - change amplitude of TWG input . 0V --> no change in baseline . 1V --> slight shift down in baseline . higher --> baseline continues to shift down . this is normal b/c of DC-balance on input to ADF - discussions w/ Dean Schamberger started to understand how this will work with real signals . important b/c baseline shifts --> inaccuracies in E-measurement o Plan - will make SPICE simulation back at Saclay about this - Dean will make measurements of DC component of BLS signal News from the Re-Baselining: ---------------------------- o Focus has been on Silicon - should hear "by the weekend" o New Trigger Costs presented (additional ~ $500K) - no real problems with this so far - people very impressed by technical progress